Another day, another few steps closer to getting the cryostat closed and pumping. We pulled the instrument out of the cryostat for the final time so that Jeff could finish installing the parts for the half-wave-plate assembly.
Jeff spent a large fraction of the day inside the cryostat. We had a little scare early in the day as Jeff was installing part of the HWP drive shaft assembly -- it got dropped on the floor, bending two of the flexible bellows in the process. After a brief moment of panic, Jeff realized he actually had spares for the parts in his parts bin, so after a quick replacement we were back on our way. After that, the wiring for the HWP system was installed and tested. The last piece of the puzzle was measuring positions of the angle encoder components on the HWP assembly so we can later infer the HWP's angle relative to the rest of the cryostat and, by extension, the sky.
I spent much of the day working on getting our optical filter, made of porous Teflon, ready to mount. This filter, which helps protect the 1 Kelvin optics from the thermal radiation from the relatively warm (~30K) rotating HWP, is a new addition for this cryostat run.
Kyle spent a lot of the day working on getting our computer network set up, which he did with the help of Milligan back in Minnesota and the CSBF networking person. He also started getting our vacuum leak checker, which arrived today from Minnesota, set up. Kyle and I both helped Jeff where we could, as he had the most difficult work for today.